0.4 L of a 8 M solution are poured into a beaker. 0.5 L of water is added to the beaker what is the new molarity?


1
Expert's answer
2021-02-10T01:38:31-0500

The initial solution contains:

n = M × V

where n - number of moles, M - molarity, V - volume.

From here:

n = 8 M × 0.4 L = 3.2 mol

The final volume of the solution is:

Vfinal = 0.4 L + 0.5 L = 0.9 L

The new molarity is:

Mfinal = n / Vfinal = 3.2 mol / 0.9 L = 3.6 M


Answer: 3.6 M